# floZ: Improved Bayesian evidence estimation from posterior samples with normalizing flows

### v2

@@ Abstract
We introduce floZ, an improved method based on normalizing flows, for estimating the Bayesian evidence (and its numerical uncertainty) from a set of samples drawn from the unnormalized posterior distribution. We validate it on distributions whose evidence is known analytically, up to 15 parameter space dimensions, and compare with two state-of-the-art techniques for estimating the evidence: nested sampling (which computes the evidence as its main target) and a $k$-nearest-neighbors technique that produces evidence estimates from posterior samples. Provided representative samples from the target posterior are available, our method is more robust to posterior distributions with sharp features, especially in higher dimensions. For a simple multivariate Gaussian, we demonstrate its accuracy for up to 200 dimensions with $10^{5}$ posterior samples. floZ has wide applicability, e.g., to estimate the evidence from variational inference, Markov Chain Monte Carlo samples, or any other method that delivers samples from the unnormalized posterior density, such as simulation-based inference. We apply floZ to compute the Bayes factor for the presence of the first overtone in the ringdown signal of the gravitational wave data of GW150914, finding good agreement with nested sampling.
@@ I Introduction
The estimation of the Bayesian evidence is in general a challenging task, as it requires averaging the likelihood over the parameters’ prior over the entire model’s parameter space. Various approaches have been proposed to this end. One that has gained particular prominence since its introduction by John Skilling Skilling2006 is nested sampling – a method designed to transform the multi-dimensional integral of the Bayesian evidence into a one-dimensional integral. Since its invention, many different algorithmic implementations of nested sampling have been proposed, which in the main are characterized by the way the likelihood-constrained step is performed: these include ellipsoidal 2009MNRAS.398.1601F , diffusive 2009arXiv0912.2380B , dynamical 2019S&C….29..891H nested sampling; nested sampling with normalizing flows Moss:2019fyi ; for a recent review, see 2023StSur..17..169B ). One of the remaining difficulties of nested sampling is its application to very large parameter spaces, where the curse of dimensionality hobbles most algorithms. Progress in this direction is being made thanks to approaches such as PolyChord 2015MNRAS.453.4384H and proximal nested sampling 2021arXiv210603646C .
@@
Other methods to compute the evidence range from the analytical (Laplace approximation and higher-order moments 10.1214/16-EJS1218 , Savage-Dickey density ratio Trotta:2005ar ) to the ones based on variations of density estimation, like parallel tempering MCMC coupled with thermodynamic integration 10.1063/1.1751356 and, more recently, on simulation-based inference with neural density (or neural ratio) estimation and/or deep learning 2023arXiv231115650K ; Jeffrey:2023stk ; 2020arXiv200410629R . A promising approach is that of evaluating the Bayesian evidence integral from a set of existing posterior samples, previously gathered e.g. via MCMC. The interest lies in the ability to obtain the evidence from the post-processing of posterior samples, which can be obtained with any suitable algorithm. Such a method, based on the distribution statistics of nearest neighbors of posterior samples, was presented in Heavens:2017afc (see also Ref. 2021arXiv211112720M for a similar approach based on harmonic reweighting of posterior samples, recently upgraded to a normalizing-flows based method in Piras2024TheFO , and Ref. Rinaldi:2024bow for a hierarchical inference approach).
@@
As first pointed out by JiaSeljak_NFEvidence_2019 , the evidence can also be extracted using normalizing flows. Building on this idea, we present an improved algorithm that provides a fast, accurate, and robust estimate of the evidence from posterior samples. Normalizing flows are a type of generative learning algorithms that aim to define a bijective transformation of a simple probability distribution into a more complex distribution by a sequence of invertible and differentiable mappings. Normalizing flows have been initially introduced in Ref. Tabak:2010 ; Tabak:2013 and then extended in various works, with applications to clustering classification Agnelli:2010 , density estimation Rippel:2013 ; Laurence:2014 , and variational inference Rezende:2015 . While normalizing flows have been used in several parameter estimation scenarios in cosmology and GW physics Green:2020hst ; Green:2020dnx ; Dax:2021tsq ; Dax:2021myb ; 2023PhRvL.130q1403D ; Wildberger:2022agw ; Bhardwaj:2023xph ; Crisostomi:2023tle ; Leyde:2023iof ; Shih:2023jme ; Vallisneri:2024xfk , here we introduce a novel training prescription that is specifically tailored towards evaluating the evidence. This paper is structured as follows: we introduce normalizing flows in Section II, and how a suitable loss can be defined in order to encode the objective of evidence estimation; in Section III, we validate our approach on tractable likelihood in up to 15 parameter space dimensions, and benchmark it against dynesty dynesty_code (an implementation of nested sampling) and the $k$-nearest neighbors method of Heavens:2017afc , demonstrating the superiority of our approach over the latter. In Section IV, we apply our method to real GW data analysis by estimating the Bayes factor in favor of the presence of the first overtone in the ringdown phase of GW150914. We conclude in Section V with an outlook on future applications.
@@ II.1 Normalizing flows
where ${\rm E}_{{\rm p}(\bm{x})}$ represents the expectation value under ${\rm p}(\bm{x})$, which can be computed by a simple average given a set of independent samples and their corresponding probability density.
@@ II.2 Evidence estimation
In the context of Bayesian analysis, ${\rm p}(\bm{x})$ represents the posterior distribution of the parameters $\bm{x}$, $\hat{\rm p}(\bm{x})$ is the product of likelihood function and prior distribution (the unnormalized posterior), and ${Z}$ is the evidence. From Eq. (3) we can write
$\frac{\hat{\rm p}(\bm{x})}{{Z}}\,\text{d}\bm{x}={\rm n}(\bm{y})\,\text{d}\bm{y}\,,$ (4)
@@
${Z}=\frac{\hat{\rm p}(\textbf{f}_{\bm{\phi}}(\bm{y}))}{{\rm n}(\bm{y})}\,\left|{\rm det}\frac{\partial\textbf{f}_{\bm{\phi}}}{\partial\bm{y}}\right|=\frac{\hat{\rm p}(\bm{x})}{{\rm n}(\textbf{f}_{\bm{\phi}}^{-1}(\bm{x}))}\,\left|{\rm det}\frac{\partial\textbf{f}_{\bm{\phi}}^{-1}}{\partial\bm{x}}\right|^{-1}\,.$ (5)
Resorting to Eq. (1), the evidence in Eq. (5) can be mapped using normalizing flows as
${Z}\mapsto\zeta(\bm{x},\bm{\phi})=\frac{\hat{\rm p}(\bm{x})}{\text{q}_{\bm{\phi}}(\bm{x})}\,.$ (6)
This computation exploits the crucial property of ${\rm n}(\bm{y})$ of being normalized by construction. Given an unnormalized statistical model $\hat{\rm p}(\bm{x})$ and a transformation $\textbf{f}_{\bm{\phi}}$, the normalization ${Z}$ can be explicitly computed as the right-hand side of Eq. (6). Note that this computation can be performed in the variable $\bm{x}$ as well as in the latent space $\bm{y}$, as shown in Eq. (5).
@@ III.3 Benchmarking
In Figure 5, we illustrate the behavior of the loss schedule for the Gaussian mixture case. The four individual loss terms are shown with different colors, the total loss in thick/black; training losses are depicted with continuous lines and validation losses with dotted ones. As described in Sec. II.4, we show that each loss term solely contributes to 20 epochs before linearly transitioning to the next term in 5 epochs. In a span of 500 epochs, the network is trained on 5 cycles of the loss terms.
@@ V Conclusions
We have introduced floZ, an improved method that uses normalizing flows for estimating the Bayesian evidence and its numerical uncertainty using pre-existing samples drawn from the target unnormalized posterior distribution, e.g. as obtained by standard MCMC methods.
@@
Pushing our approach to much larger parameter spaces remains challenging because of two limitations, both fundamentally stemming from the curse of dimensionality: first, obtaining posterior samples in high dimensions becomes more difficult (although some methods, like Gibbs sampling and HMC, can show mild scaling with dimensionality in some favorable circumstances); second, training the flow in higher dimensions suffers from increasing inaccuracies (but see Piras2024TheFO for a normalizing-flows-enhanced harmonic mean estimator with similar performance as our own up to 150 dimensions, at least for the restricted case of unimodal, mildly non-Gaussian targets). However, we note that we only need accurate densities in latent space near the peak (as opposed to into the tails of the distribution), which opens the door to potentially exploiting the existence of a well-defined ‘typical set’ in high dimensions, a consequence of the concentration of measure phenomenon. We shall explore this idea further in a dedicated future paper.
@@
Data availability: The floZ code and all the necessary files to reproduce the results in this paper will soon be made available.

### v3

# Bayesian evidence estimation from posterior samples with normalizing flows

@@ Abstract
We propose a novel method (floZ), based on normalizing flows, to estimate the Bayesian evidence (and its numerical uncertainty) from a pre-existing set of samples drawn from the unnormalized posterior distribution. We validate it on distributions whose evidence is known analytically, up to 15 parameter space dimensions, and compare with two state-of-the-art techniques for estimating the evidence: nested sampling (which computes the evidence as its main target) and a $k$-nearest-neighbors technique that produces evidence estimates from posterior samples. Provided representative samples from the target posterior are available, our method is more robust to posterior distributions with sharp features, especially in higher dimensions. For a simple multivariate Gaussian, we demonstrate its accuracy for up to 200 dimensions with $10^{5}$ posterior samples. floZ has wide applicability, e.g., to estimate evidence from variational inference, Markov Chain Monte Carlo samples, or any other method that delivers samples and their likelihood from the unnormalized posterior density. As a physical application, we use floZ to compute the Bayes factor for the presence of the first overtone in the ringdown signal of the gravitational wave data of GW150914, finding good agreement with nested sampling.
@@ I Introduction
The estimation of the Bayesian evidence is generally a challenging task, as it requires averaging the likelihood over the prior density in the model’s parameter space. Various approaches have been proposed to this end. One that has gained particular prominence since its introduction by John Skilling Skilling2006 is nested sampling – a method designed to transform the multi-dimensional integral of the Bayesian evidence into a one-dimensional integral. Since its invention, many different algorithmic implementations of nested sampling have been proposed, which in the main are characterized by the way the likelihood-constrained step is performed: these include ellipsoidal 2009MNRAS.398.1601F , diffusive 2009arXiv0912.2380B , dynamical 2019S&C….29..891H nested sampling; nested sampling with normalizing flows Moss:2019fyi ; Nessai_2021 ; for a recent review, see 2023StSur..17..169B ). One of the remaining difficulties of nested sampling is its application to very large parameter spaces, where the curse of dimensionality hobbles most algorithms. Progress in this direction is being made thanks to approaches such as PolyChord 2015MNRAS.453.4384H and proximal nested sampling 2021arXiv210603646C .
@@
Other methods to compute the evidence range from the analytical (Laplace approximation and higher-order moments 10.1214/16-EJS1218 , Savage-Dickey density ratio Trotta:2005ar ) to the ones based on variations of density estimation, like parallel tempering MCMC coupled with thermodynamic integration 10.1063/1.1751356 and, more recently, on simulation-based inference with neural density (or neural ratio) estimation and/or deep learning 2023arXiv231115650K ; Jeffrey:2023stk ; 2020arXiv200410629R . Ref. Mancini+2023 discussed evaluating the evidence using a normalizing flow and the ratio of a neural posterior estimator and neural likelihood estimator. However, the ratio of the two approximate quantities resulted in a larger compound error in the evidence.
Another particularly interesting method was proposed in JiaSeljak_NFEvidence_2019 , where Optimal Bridge Sampling Meng1996 ; chen2012monte is improved by using normalizing flows to obtain a suitable density estimator of the proposal distribution which has large overlap with the posterior.
A promising approach is that of evaluating the Bayesian evidence integral from a set of existing posterior samples, previously gathered e.g. via MCMC. The interest lies in the ability to obtain the evidence from the post-processing of posterior samples, which can be obtained with any suitable algorithm. Such a method, based on the distribution statistics of nearest neighbors of posterior samples, was presented in Heavens:2017afc (see also Ref. 2021arXiv211112720M for a similar approach based on harmonic reweighting of posterior samples, recently upgraded to a normalizing-flows based method in Polanska:2024arc ; Piras2024TheFO , and Ref. Rinaldi:2024bow for a hierarchical inference approach). Here, we propose a new method based on normalizing flows.
@@
Normalizing flows are a type of generative learning algorithms that aim to define a bijective transformation of a simple probability distribution into a more complex distribution by a sequence of invertible and differentiable mappings. Normalizing flows have been initially introduced in Ref. Tabak:2010 ; Tabak:2013 and then extended in various works, with applications to clustering classification Agnelli:2010 , density estimation Rippel:2013 ; Laurence:2014 , and variational inference Rezende:2015 . While normalizing flows have been used in several parameter estimation scenarios in cosmology and GW physics Green:2020hst ; Green:2020dnx ; Dax:2021tsq ; Dax:2021myb ; 2023PhRvL.130q1403D ; Wildberger:2022agw ; Bhardwaj:2023xph ; Crisostomi:2023tle ; Leyde:2023iof ; Shih:2023jme ; Vallisneri:2024xfk , here we introduce them, together with a novel training prescription, as a method to evaluate the evidence. This paper is structured as follows: we introduce normalizing flows in Section II, and how a suitable loss can be defined in order to encode the objective of evidence estimation; in Section III, we validate our approach on tractable likelihood in up to 15 parameter space dimensions, and benchmark it against dynesty dynesty_code (an implementation of nested sampling) and the $k$-nearest neighbors method of Heavens:2017afc , demonstrating the superiority of our approach over the latter. In Section IV, we apply our method to real GW data analysis by estimating the Bayes factor in favor of the presence of the first overtone in the ringdown phase of GW150914. We conclude in Section V with an outlook on future applications.
@@ II.1 Normalizing flows
where ${\rm E}_{{\rm p}(\bm{x})}$ represents the expectation value under ${\rm p}(\bm{x})$, which can be computed by a simple average given a set of independent samples.
@@ II.2 Evidence estimation
In the context of Bayesian analysis, ${\rm p}(\bm{x})$ represents the posterior distribution of the parameters $\bm{x}$, $\hat{\rm p}(\bm{x})$ is the product of likelihood function and prior distribution (the unnormalized posterior), and ${Z}$ is the evidence.
From Eq. (1), the evidence in Eq. (3) can be written as
${Z}\mapsto\zeta(\bm{x},\bm{\phi})=\frac{\hat{\rm p}(\bm{x})}{\text{q}_{\bm{\phi}}(\bm{x})}\,,$ (4)
@@
${Z}\mapsto\frac{\hat{\rm p}(\textbf{f}_{\bm{\phi}}(\bm{y}))}{{\rm n}(\bm{y})}\,\left|{\rm det}\frac{\partial\textbf{f}_{\bm{\phi}}}{\partial\bm{y}}\right|=\frac{\hat{\rm p}(\bm{x})}{{\rm n}(\textbf{f}_{\bm{\phi}}^{-1}(\bm{x}))}\,\left|{\rm det}\frac{\partial\textbf{f}_{\bm{\phi}}^{-1}}{\partial\bm{x}}\right|^{-1}\,.$ (5)
This computation exploits the crucial property of ${\rm n}(\bm{y})$ of being normalized by construction. In this paper, we evaluate $\hat{\rm p}(\bm{x})$ by multiplying the value of the likelihood function (saved in the output at each step of an MCMC chain) by the prior density at the same location. We note however that our method is generally applicable to any collection of posterior samples for which access to the unnormalized posterior $\hat{\rm p}(\bm{x})$ is available. For example, in the context of simulation-based inference, one could use neural conditional density estimation to learn the likelihood function (doi:10.1073/pnas.1912789117, ), then perform traditional MCMC with the neural likelihood to obtain samples from the posterior.
Given an unnormalized statistical model $\hat{\rm p}(\bm{x})$ and a transformation $\textbf{f}_{\bm{\phi}}$, the normalization ${Z}$ can be explicitly computed as the right-hand side of Eq. (4). Note that this computation can be performed in the variable $\bm{x}$ as well as in the latent space $\bm{y}$, as shown in Eq. (5).
@@ III.3 Benchmarking
Figure 4 compares the evidence estimate from $floZ$, $kNN$, and nested sampling for the four 2-dimensional posterior distributions. In Figure 5, we illustrate the behavior of the loss schedule for the Gaussian mixture case. The four individual loss terms are shown with different colors, the total loss in thick/black; training losses are depicted with continuous lines and validation losses with dotted ones. As described in Sec. II.4, we show that each loss term solely contributes to 20 epochs before linearly transitioning to the next term in 5 epochs. In a span of 500 epochs, the network is trained on 5 cycles of the loss terms.
@@ V Conclusions
We have introduced floZ, a novel method that uses normalizing flows for estimating the Bayesian evidence and its numerical uncertainty using pre-existing samples drawn from the target unnormalized posterior distribution, e.g. as obtained by standard MCMC methods.
@@
Pushing our approach to much larger parameter spaces remains challenging because of two limitations, both fundamentally stemming from the curse of dimensionality: first, obtaining posterior samples in high dimensions becomes more difficult (although some methods, like Gibbs sampling and HMC, can show mild scaling with dimensionality in some favorable circumstances); second, training the flow in higher dimensions suffers from increasing inaccuracies (but see Piras2024TheFO for a normalizing-flows-enhanced harmonic mean estimator with similar performance as our own up to 150 dimensions, at least for the restricted case of unimodal, mildly non-Gaussian targets). However, we note that we only need accurate densities in latent space near the peak (as opposed to into the tails of the distribution), which opens the door to potentially exploiting the existence of a well-defined ‘typical set’ in high dimensions, a consequence of the concentration of measure phenomenon. An interesting possibility for improvement over the aforementioned inaccuracies of the trained flow, potentially allowing for enhanced precision, is provided by importance sampling 2023PhRvL.130q1403D , where importance weights are applied to correct the distribution of the estimated evidence, based on the underlying posterior distribution and its recovery with the flow. We shall explore these ideas further in a dedicated future paper.
@@
Data availability: The floZ code to reproduce the results in this paper, and to estimate the evidence for any set of posterior samples, likelihood, and prior densities is available at https://github.com/Rahul-Srinivasan/floZ.